# Before & After: The Effect of EU’s 2022 Code of Practice on Disinformation

### v1

@@ 6. Served Advertisements
Table 2. Ad network operators serving programmatic ads to misinformation websites. Even though Google and Microsoft have signed the Code of Practice on Disinformation, they still deliver ads, and therefore revenue, to misinformation websites.
6.1. Ad Networks
Using data coming from 6 distinct user profiles (2 genders and 3 countries) and one clean user profile, we detect ads from 2,480 advertisers in 392 misinformation websites. We attribute the large number of distinct advertisers to the fact that we use different vantage points (i.e., 3 geographic locations), thus being served ads from local brands.
First, we study the ad networks that deliver ads to misinformation websites. We achieve this by studying the ad URLs the crawler detected in any of the used personas. We analyze these ad URLs and extract the domain from them. Then, we group domains on the eTLD+1 level. For example, ads served by adclick.g.doubleclick.net and googleads.g.doubleclick.net are all grouped under the doubleclick.net domain. Finally, to understand the actual entities involved in serving these ads, we group domains based on the companies that own these domains. For example, ads served by doubleclick.net and googleadservices.com are all attributed to Google LLC. To identify the entities that own and operate each domain, we utilize the DuckDuckGo Tracker Radar dataset (trackerRadar, ). We acknowledge that there might be various intermediaries involved in the (re)selling of ad inventory but this analysis is outside the scope of this work. We summarize our findings in Table 2.
@@ 6.1. Ad Networks
We find that there are a lot of legitimate and popular ad networks that still deliver ads to misinformation websites. We find that Google is the most prominent ad facilitator of misinformation websites serving ads to 69% of the misinformation websites, which contained ads. The second most popular ad network that serves ads to misinformation websites is RevContent. Previous work has already demonstrated that RevContent is popular among misinformation websites (papadogiannakis2023funds, ; han2022infrastructure, ). Unfortunately, we find that ads are served on misinformation websites even by entities that have signed the Code of Practice on Disinformation (e.g., Google and Microsoft) (copodSignatories, ), suggesting that there is still work to be done.
@@
It is important to mention that these findings highlight even more the issue described in previous sections with the analysis of ads.txt and sellers.json files. An ad network could argue that they have indeed authorized some misinformation websites to enter their platform (i.e., findings of ads.txt and sellers.json analysis) but they have now blocked them and no longer work with them. However, the findings of this analysis are definitive proof that some ad networks still have business relationships with misinformation websites. These ad networks serve the actual ads and eventually pay money to publishers of misinformation content.
@@
To make matters worse, we discover that there exist ad networks that deliver ads of misinformation websites to users. For example, when our female USA-based persona visited the website theconservativebried.com, Google served an ad for goop.com, a website associated with pseudoscience and dozens of failed fact-checks (mbfcGoop, ). We observe similar behavior for various ad networks, including Google, Outbrain, RevContent and Yandex LLC. Evidently, ad networks do not properly review the content they promote, thus disseminating misinformation. It is important to notice that the CoP explicitly states that signatories should work towards stopping ads that contain false claims (Commitment 2: “TACKLING ADVERTISING CONTAINING DISINFORMATION”) and Google is one of the signatories. All the above suggest that there even though the CoP is a step towards the right direction, more effort is required to tackle the monetization and propagation of online misinformation.
@@
Finding 4: We discover that the most popular ad networks do not closely examine their ad placements. They still still serve ads, and thus ad revenue, to almost 400 misinformation websites, and these ads might even promote false claims. This is true even for Google and Microsoft, signatories of the CoP, that place ads in 70% and 9% (resp.) of misinformation websites in our dataset.
@@ 6.2. Brand Safety of Advertisers
Next, we analyze the actual advertisers that appear next to misinformation content. We utilize the dataset of the previous section but now focus on the landing pages users arrive when they click on ads next to misinformation content. We find that there exist a lot of advertisers whose ads are placed in dozens of different misinformation websites. This behavior is consistent for various types of ads and for all personas. For example, using the USA-based female persona, we find that jjshouse.com, a popular online store for wedding dresses was advertised in 80 distinct misinformation websites. Similarly, the Greece-based male persona was served ads from Temu in 86 distinct misinformation websites (22%) while ads from overnightglasses.com were served to the male US-based user in 59 distinct misinformation websites. We should note that this finding is not a finger pointing exercise for these advertisers and does not indicate that these advertisers choose to appear next to misinformation content. Most likely, these ads are placed next to misinformation because of the ad network believing they are of interest to the virtual user. However, we should note that this can have a big impact on advertisers (i.e., brand safety) if visitors associate misinformation content with specific services or products.
Finally, we analyze the significance of the advertisers found in misinformation websites. To our surprise, we find ads from very reputable companies appearing next to misinformation content. We provide screenshots of some examples (openSourceData, ). To make matters worse, we find 23 Fortune 500 companies advertising in misinformation websites, threatening their reputation and consumer trust due to negative brand associations.
Finding 5: Ads from very popular and renowned brands (even from 23 of the largest companies in the world) are served next to misinformation content. Users can see an ad of a specific brand in up to 86 distinct misinformation websites.
@@ 7. Related Work
In (bozarth2021market, ), authors studied fake news and low-quality news websites and found that fake news publishers monetize their content using reliable ad servers. In (Broniatowski2023, ), authors studied the monetization schemes of misinformation websites, focusing on anti-vaccine pages, and observed that such websites mainly use ads to monetize their content. In (Thirumuruganathan2021, ), authors performed a cost-based analysis on fake news interventions and proposed a model that can help combat the proliferation of misinformation websites when there is uncertainty if the target indeed serves misinformation. In (papadopoulos2023fndaas, ), authors proposed a content-agnostic fake news detection service using network and structural characteristics, including requests towards third-party advertising services. A similar approach was followed in (hounsel2020identifying, ), where authors utilized non-perceptual features (e.g., domain name, DNS config) to train a multi-class model that detects disinformation websites in the wild. Their system was deployed on over 1.3M websites and was able to discover two unlisted disinformation websites. In (chalkiadakis2021rise, ), authors studied the network traffic of fake news websites and found that fake news websites are mainly accessed directly and that some popular third-party services have higher presence in fake news sites compared to real news. In (pachilakis2021youradvalue, ), authors compared advertising prices for desktop and mobile users and showed how these prices have grown over the years.
@@
The ads.txt standard was introduced back in 2017 and in (bashir2019longitudinal, ), authors performed a longitudinal study of ads.txt files and their adoption. Similar to our work, in (papadogiannakis2023funds, ), authors fetched and processed ads.txt and sellers.json files to study the business relationships between fake news websites and ad networks. They discovered that fake news websites rely on popular and credible ad networks to generate revenue. In (vekaria2023inventory, ), authors measured the compliance of the ad ecosystem with the ads.txt and sellers.json standards. Additionally, they explored how publishers use ads.txt files to form dark pools, a way to circumvent ad network restrictions to monetize misinformation content. Similarly, in (papadogiannakis2023pooling, ), authors further explored dark pooling along with hidden intermediaries and highlighted the importance of limiting the advertising revenue generated by misinformation content publishers. In (vekaria2024turning, ), authors perform an extensive study and demonstrated that dark pooling can be diminished by notifying stakeholders, and most importantly ad networks.
Previous work has also focused on advertisements and monetization of misinformation websites in the past. In (han2022infrastructure, ), authors discovered that ad networks such as RevContent, Google DoubleClick and Outbrain disproportionately support misinformation websites. In addition to this, they demonstrated how misinformation websites rely on legitimate infrastructure providers (e.g., Cloudflare, GoDaddy) to support misinformation dissemination. In (zeng2020bad, ), author studied the content of ads served on either real or fake news websites and found that almost half of the ads they labeled were problematic. In (broniatowski2023measuring, ), authors compared pro-vaccine and anti-vaccine websites and found that pro-vaccine follow more aggressive monetization strategies. Taboola was the ad networks that served the most problematic ads. In (ahmad2024companies, ), authors demonstrated that ads largely finance misinformation websites and that can have a great effect on advertisers’ brand safety. In (bellman2018brand, ), authors performed a user study to understand how questionable content affects the brand of pre-roll advertisers. Similarly, in (shehu2021risk, ), the authors studied the effect of website quality on advertisers that utilize programmatic advertising. In (braun2019fake, ), authors interviewed actual stakeholders involved in online advertising and explain that some advertisers might tolerate fake news websites because they can reach users at a lower cost. In (10.1007/978-3-030-50232-4_6, ), authors created a misinformation website and disseminated fake news articles as ads in Facebook to study how different generations interact with misinformation. Finally, previous work has studied disinformation campaigns that utilized Facebook ads to affect ellections (cano2021disinformation, ; alvarez2020good, ).
@@ 8.1. Discussion
The Strengthened Code of Practice on Disinformation is a step towards the right direction, but there is still work to be done. The goal of this work is not to discredit the effort of the CoP but rather to highlight the importance of scrutinizing ad placements and marking more misinformation websites as ineligible for monetizaion. We believe that ad networks should work towards ensuring that misinformation is no longer profitable and that even popular misinformation websites can no longer survive because of ad revenue. Reducing the financial incentives of misinformation spreaders can increase the quality of information and trust in factual reporting. Finally, the self-regulatory nature of the CoP is a positive indication of desire for a more reliable news ecosystem but unfortunately its adoption is limited, with only 44 signatories of various categories (e.g., ad networks, fact checkers, social media platforms) as of July 2024. It is evident that current signatories have been laying the foundation, but we hope new signatories from the online ad ecosystem (i.e., ad networks) will take initiative for more substantial actions.
It is customary in such studies to have a control group - that is, a set of non-disinformation websites and study what is the change in advertisements in these websites as well. Such a control group enables researchers to argue whether a change they see in misinformation websites is due to the CoP or whether this change happens in non-misinformation websites as well. However, this is not the case for our study. Indeed, we are not interested in attributing any changes we see to the CoP (or not). The goal of this work is to discover We just want to see if the change that the CoP expected (i.e., that disinformation websites will not have ads) happened and to what extent. To make this conclusion we do not need a control group.
8.2. Ethical Considerations
In this work, we make concentrated effort to not affect the performance of any Web service in any way. To download ads.txt and sellers.json files, we utilize existing crawlers and reach each domain only once to download the specific file. This behavior is far less intrusive compared to the automated processes that DSPs and SSPs use daily to verify programmatic advertising. Additionally, our data was collected once and all of our analysis is performed offline. Regarding the collection of ads, we ensure that we click on each ad URL only once, similarly to methodologies presented in previous work (papadogiannakis2023funds, ; vekaria2024turning, ). Since previous work has demonstrated that the cost per thousand ad impressions is only a few dollars (pachilakis2021youradvalue, ), we argue that our study’s influence on the advertising ecosystem is negligible. Finally, following the GDPR and ePrivacy regulations, we do not collect or process any personal information of real users.
@@ 8.3. Conclusion
The Strengthened Code of Practice on Disinformation was signed in June 2022, in an attempt to fight disinformation in a self-regulatory manner. The first commitment of the CoP is that entities involved in ad placements demonetize dissemination of disinformation. The expectation is that since the CoP was signed by leading industry players (e.g., Google and Microsoft), two years after the introduction of the Strengthened CoP, things will have improved. In this work, we perform a historical analysis and demonstrate that this is not the case. In fact, misinformation websites are still able to retain business contracts with popular ad networks and monetize their content. We show that recent regulation, and adoption of Code of Practice for Disinformation has not made a significant impact on the misinformation ecosystem, and that ad networks stopped working with only insignificant misinformation websites that do not have a lot of visitors, i.e., traffic. Additionally, we discover that signatories of the CoP do not differentiate between unreliable and reliable news sources. Finally, we perform a dynamic analysis of ads served on misinformation websites. We establish that not only popular ad networks still facilitate the monetization of the majority of misinformation websites, but that they also serve ads of Fortune 500 companies next to misinformation content, threatening the companies’ reputation.

### v2

@@ 6.1. Ad Networks
Table 2. Ad network operators serving programmatic ads to misinformation websites. Even though Google and Microsoft have signed the CoP, they still deliver ads, and therefore revenue, to misinformation websites.
@@
We find that there are a lot of legitimate and popular ad networks that still deliver ads to misinformation websites. Google is the most prominent ad facilitator of misinformation websites, serving ads to 69% of the misinformation websites that contained ads. The second most popular ad network that serves ads to misinformation websites is RevContent, known to be popular among misinformation websites (Papadogiannakis et al., 2023; Han et al., 2022). Unfortunately, we find that ads are served on misinformation websites even by entities that have signed the Code of Practice on Disinformation (e.g., Google and Microsoft) (Comission, 2022b), suggesting that there is still work to be done.
@@
It is important to mention that these findings highlight the issue described in previous sections with the analysis of ads.txt and sellers.json files. Ad networks may claim they once allowed misinformation websites (i.e., findings of previous sections) but later blocked them. Our analysis proves some still partner with misinformation publishers, serving ads and funding them.
@@
To make matters worse, we discover ad networks serving misinformation ads. For example, when our USA-based female persona visited the website theconservativebried.com, Google served an ad for goop.com, a website linked to pseudoscience and failed fact-check (Check, 2024a). We observe similar behavior with Outbrain, RevContent and Yandex LLC. Evidently, ad networks do not properly review the content they promote, thus disseminating misinformation. The CoP explicitly states that signatories (such as Google) should work towards stopping ads that contain false claims (Commitment 2: “TACKLING ADVERTISING CONTAINING DISINFORMATION”). All the above suggest that even though the CoP is a step towards the right direction, more effort is required to tackle the monetization and propagation of online misinformation.
@@
Finding 4: We discover that the most popular ad networks do not closely examine their ad placements. They still serve ads, and thus ad revenue, to almost 400 misinformation websites. Google and Microsoft, signatories of the CoP, place ads in 70% and 9% (resp.) of misinformation websites in our dataset.
@@ 6.2. Brand Safety of Advertisers
Next, we analyze the actual advertisers that appear within the misinformation context. We utilize the dataset of the previous section, but now focus on the landing pages users arrive when they click on ads displayed on misinformation websites. We find that there exist a lot of advertisers whose ads are placed in dozens of different misinformation websites. This behavior is consistent for various types of ads and for all personas. For example, using the USA-based female persona, we find that jjshouse.com, a popular online store for wedding dresses, was advertised in 80 distinct misinformation websites. Similarly, the Greece-based male persona was served ads from Temu in 86 distinct misinformation websites (22%) while ads from overnightglasses.com were served to the male US-based user in 59 distinct misinformation websites. We should note that this finding is not a finger pointing exercise for these advertisers and does not indicate that these advertisers choose to appear in misinformation websites. Most likely, these ads are placed next to misinformation because of the ad network believing they are of interest to the virtual user. However, we should note that this can have a big impact on advertisers (i.e., brand safety) if visitors associate misinformation content with specific services or products.
Finally, we analyze the significance of the advertisers found in misinformation websites. To our surprise, we find ads from very reputable companies appearing in misinformation websites. We provide screenshots of examples (Papadogiannakis, 2024). We find 23 Fortune 500 companies advertising in misinformation websites, threatening their reputation and consumer trust due to negative brand associations.
Finding 5: Ads from very popular and renowned brands (even from 23 of the largest companies in the world) are served in misinformation websites. Users can see an ad of a specific brand in up to 86 distinct misinformation websites.
@@ 7. Related Work
The monetization of misinformation content through online ads has been thoroughly studied. In fact, in (Thirumuruganathan et al., 2021), authors performed a cost-based analysis on fake news interventions and proposed a model that can help combat the proliferation of misinformation websites. In (Bozarth and Budak, 2021), authors studied fake news and low-quality news websites and found that fake news publishers monetize their content using reliable ad servers. Similarly, in (Broniatowski et al., 2023), authors studied the monetization schemes of misinformation websites, focusing on anti-vaccine pages, demonstrating they mainly rely on ads. In (Chalkiadakis et al., 2021), authors found that some popular third-party services have higher presence in fake news sites compared to real news. In (Han et al., 2022), expanded upon this finding and found that misinformation websites rely on popular Web services such as Cloudflare, Google and RevContent to host and monetize their content. This behavior is leveraged by (Papadopoulos et al., 2023), where authors proposed a content-agnostic fake news detection service using network and structural characteristics (including requests towards third-party advertising services), and (Hounsel et al., 2020), authors utilized non-perceptual features (e.g., domain name, DNS config) to train a multi-class model that detects disinformation websites.
@@
Regarding the ads.txt and sellers.json standards that are the foundation of our methodology, in (Bashir et al., 2019), authors performed a longitudinal study of ads.txt files and their adoption. In (Vekaria et al., 2024a), authors also measured the compliance of the ad ecosystem with ad standards and explored how publishers use ads.txt files to form dark pools, a way to circumvent restrictions and monetize misinformation content. Similarly, in (Papadogiannakis et al., 2025), authors further explored dark pooling along with hidden intermediaries and highlighted the importance of limiting the ad revenue generated by misinformation publishers. In (Vekaria et al., 2024b), authors demonstrated that dark pooling can be diminished by notifying stakeholders and ad networks. In this work, we expand the methodology and make use of the dataset presented in (Papadogiannakis et al., 2023), where authors processed ads.txt and sellers.json files to study the business relationships between fake news websites and ad networks. They discovered that fake news websites rely on popular and credible ad networks to generate revenue.
Previous work has examined the content of ads served on misinformation websites, finding that nearly half of the ads were problematic (Zeng et al., 2020) and that some advertisers tolerate fake news websites due to lower costs (Braun and Eklund, 2019). In this work, we mainly focus on the ad networks responsible for serving ads on misinformation websites and only briefly examine the advertisers to discuss brand safety. Brand safety has been more extensively studied in (Ahmad et al., 2024), where authors demonstrated that ads largely finance misinformation websites and that can have a great effect on advertisers’ brand safety. Previous work has also studied how questionable content affects the brand of pre-roll advertisers (Bellman et al., 2018) and advertisers that utilize programmatic advertising (Shehu et al., 2021), and how reputable brands can appear next to illicit content, unbeknownst to them (Papadogiannakis et al., 2024).
@@ 8.1. Discussion
The Strengthened Code of Practice on Disinformation is a step towards the right direction, but more work is needed. This study doesn’t aim to discredit the CoP but to stress the need for better scrutiny of ad placements and more misinformation websites marked ineligible for monetization. We believe ad networks must ensure misinformation is no longer profitable, preventing even popular websites from surviving on ad revenue. Reducing financial incentives for misinformation can improve information quality and trust in factual reporting. Although the self-regulatory nature of the CoP is promising, its adoption is limited, with only 44 signatories as of July 2024. It is evident that current signatories have been laying the foundation, but we hope new signatories from the online ad ecosystem (i.e., ad networks) will take initiative for more substantial actions.
It is customary in such studies to have a control group, that is, a set of non-disinformation websites and study what is the change in advertisements in these websites as well. Such a control group enables researchers to argue whether a change they see in misinformation websites is due to the CoP. However, this is not needed for this work. Indeed, we are not interested in attributing any changes we see to the CoP (or not). The goal of this work is to discover if the change that the CoP expected (i.e., that disinformation websites will not have ads) happened and to what extent.
Finally, we acknowledge the argument that some ad networks might still work with misinformation websites because they classify these websites as serving “low credibility content” rather than deliberate disinformation. Classifying a website as misinformation or not falls outside the scope of this work. We rely on external sources (e.g., MBFC, ScienceFeedback) that work on this exact topic and the academic community has already reviewed their reliability.
@@ 8.2. Conclusion
The Strengthened Code of Practice on Disinformation was signed in June 2022, in an attempt to fight disinformation in a self-regulatory manner. The first commitment of the CoP is that entities involved in ad placements demonetize dissemination of disinformation. The expectation is that since the CoP was signed by leading industry players (e.g., Google and Microsoft), two years after the introduction of the strengthened CoP, things will have improved. In this work, we perform a historical analysis and demonstrate that this is not the case. In fact, misinformation websites are still able to retain business contracts with popular ad networks and monetize their content. We show that the adoption of the Code of Practice for Disinformation has not made a significant impact on the misinformation ecosystem, and that ad networks stopped working with only insignificant misinformation websites that do not have a lot of visitors. Finally, we perform a dynamic analysis of ads served on misinformation websites. We establish that not only popular ad networks still facilitate the monetization of the majority of misinformation websites, but that they also serve ads of respectable and popular brands next to misinformation content.
Acknowledgements.
Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or European Research Executive Agency (REA). Neither the European Union nor the granting authority can be held responsible for them.
Appendix A User Profiles
We form browser profiles to emulate real and distinct Web users. Our methodology revolves around the idea that user profiles can be formed based on websites that are representative of the user’s gender (Agarwal et al., 2020). For example, if websites a.com, b.com and c.com are mainly visited by female users from the Netherlands, and we visit these websites with a Dutch IP address using a clean user profile, third parties will deduce that the profile belongs to a female user from the Netherlands.
To form a list of websites that are associated with a gender of a specific country, we utilize the SimilarWeb analytics platform (LTD, 2023). Specifically, we visit SimilarWeb and extract the 5 most popular websites in each category and each country of our experiment. SimilarWeb uses a taxonomy system with 210 distinct website categories (including subcategories). For each website, we extract its demographic data (if available). SimilarWeb uses two gender groups (i.e., female and male) to classify visitors of websites. Next, we sort websites based on their audience’s gender representation and ensure that all websites have a representation of at least 55% for that gender. We set an upper threshold of 200 websites and a lower threshold of 100 websites to train each persona. The lower threshold is more than enough, since previous work has demonstrated that a persona is stable after visiting the top 10 websites associated with users of that demographic (Agarwal et al., 2020). This process leads to 6 (combination of 2 genders and 3 countries) lists of websites that represent the interests of the specific persona.
We utilize Playwright (Microsoft, 2024), a browser automation framework and set up legitimate browser profiles, retaining cookies and local storage. To build the user profiles, we try to emulate the behavior of a real user that visits some websites to consume online content. We try to make the behavior of our automated crawler as non-deterministic as possible to appear as a legitimate user. The behavior we are emulating is that of a user that visits a website, reads its content, and then navigates through the website after quickly inspecting the content of each page. For each website in our target list, we (i) visit the landing page, (ii) wait for a random number of seconds in range [2, 5], and (iii) explore 5 subpages. In order to explore a subpage, our crawler clicks on a randomly selected internal URL and navigates to this subpage, waits for a random number of seconds to emulate a user consuming content, scrolls for a random number of seconds and then waits for an additional random number of seconds. The websites of the list are visited in a random order and after all websites have been visited, the user profile is stored for further use.
Appendix B Popularity Evolution
We find that popular ad networks no longer work with websites with low traffic. The classification of a website’s popularity is based on traffic data from 2024. We study if their popularity was similar when the ad networks made their decisions. We compare the popularity rank of the 461 misinformation websites of Figure 3. We make use of the Tranco list and compare their ranks during June 2021 and June 2023 (i.e., one year before and one year after the CoP was signed). We find that for the websites in both lists, their rank distribution is somewhat similar. In fact, the mean rank decreased by -9% (i.e., more popular) during this 2-year period, while the median rank increased 42%.
Next, we perform a 2-sample KS test for the ranks of misinformation websites in June 2021 and June 2023 and find that the p-value is large enough (i.e., 0.16) so we cannot reject the hypothesis that the two distributions are the same. In addition to this, we argue that the difference in the median rank might be credited to some extent the variance the Tranco list exhibits from one version to the other. We perform exactly the same experiment for the same time period for the real news websites published in (Papadogiannakis et al., 2023). We find that for real news websites, the median rank increased by 35%, a comparable magnitude. Altogether, we demonstrate that the popularity of the misinformation websites we studied has not changed significantly before and after the CoP was signed, thus our argument that ad networks have dropped only unpopular misinformation websites still holds. We selectively compare the rank of misinformation websites in June 2023 and not January 2024 (just like we do in Figure 3) because Tranco has since changed the ground truth datasets they utilize to compute the aggregated ranks.
Appendix C Popular Misinformation Websites
To our surprise, we find that even today, there are a lot of ad networks working with certain misinformation websites that one can hardly argue they are legitimate. For example, we find that there are at least 10 of the most popular ad networks (including Google, Amazon, AppNexus and Pubmatic) that have an explicit and direct business relationship with worldlifestyle.com. According to MediaBias/FactCheck, WordLifestyle is “a strong pseudoscience website based on the promotion of false or misleading claims regarding science and health” and it has a low factual reporting score “due to a lack of sourcing and blatant clickbait abuse to generate advertising revenue” (Check, 2024f). Even such websites that are known to take advantage of the ad ecosystem in order to monetize misinformation still have direct business contracts with popular ad networks. In fact, by manually visiting this website, we find that Google indeed serves ads to visitors.
In addition to this, we find that the website of “One America News Network” (OANN) is an approved ad inventory seller within the most popular ad networks (including Google and AppNexus, signatories of the Strengthened Code of Practice on Disinformation). OANN is considered a far-right biased news website that promotes propaganda, conspiracy theories and has over 10 failed fact checks (Check, 2024c). According to SimilarWeb, this website achieves over 2.5M monthly visits, and in conjunction with the fact that almost 95% of the visitors come from the United States, an advertising valuable country, it can generate substantial ad revenue.
Appendix D Active Misinformation Websites
Inspired by previous findings that misinformation websites might have a short longevity (Chalkiadakis et al., 2021), we aim to refine our list by accumulating only misinformation websites which are still active. Towards this extent, we perform a manual analysis using independent reviewers. We start from the extended list of 2,469 misinformation websites and using an automated crawler we take screenshots of the entire landing page of each website. Our crawler was able to successfully access 2,049 websites and capture screenshots. Next, two independent reviewers coming from a computer science background (not the authors of this work) evaluate which websites are still active based on the content of the landing pages. Reviewers are allowed to classify each website as either “Active” or “Inactive”. Moreover, reviewers are allowed to manually visit a website and navigate through its pages if they are not certain which label to assign simply based on the screenshot of the landing page. The two reviewers are independent of each other and completed the task in a different environment, without influencing one another.
Both reviewers evaluated the entire list of 2,469 misinformation websites and agreed to include 1,889 websites, agreed to exclude 568 websites and disagreed on 11 websites. The reviewers agreed in 99.55% of websites and had a Cohen Kappa score of 0.99. Note that a high inter-rater agreement score is expected in this context because deciding if a website is active is an easy task for a human evaluator. We decide to only accept websites that both raters agree are active and end up with a list of 1,889 active misinformation websites. We make this list of active misinformation websites publicly available (Papadogiannakis, 2024).
Appendix E Ethical Considerations
In this work, we make concentrated effort to not affect the performance of any Web service in any way. To download ads.txt and sellers.json files, we utilize existing crawlers and reach each domain only once to download the specific file. This behavior is far less intrusive compared to the automated processes that DSPs and SSPs use daily to verify programmatic advertising. Additionally, our data was collected once and all of our analysis is performed offline. Regarding the collection of ads, we ensure that we click on each ad URL only once, similarly to methodologies presented in previous work (Papadogiannakis et al., 2023; Vekaria et al., 2024b). Since previous work has demonstrated that the cost per thousand ad impressions is only a few dollars (Pachilakis et al., 2022), we argue that our study’s influence on the advertising ecosystem is negligible. Finally, following the GDPR and ePrivacy regulations, we do not collect or process any personal information of real users.